# Addressing Diverging Training Costs using Local Restoration for Precise Bird’s Eye View Map Construction

### v3

@@ IV-C Manipulation of BEV Scope and Resolution
Figure 5: Comparison on increasing costs to BEVFusion. We incrementally add an MSA layer to the fusion neck and compare their costs. X and Y axes denote costs and the number of MSA layers, respectively.
(a) Green ROIs: Far regions occluded by cars.
(b) Green ROIs: Nearby Regions Occluded by White and Green Trucks, Brown ROIs: Faded Far regions.
Figure 6: Overconfident Issues. Our map construction struggles with the overconfident prediction of uncertain regions.
Example. Our ×$\times 4$ TNN employs (-52.0m, 52.0m, 2.0m/px) and (-50m, 50m, 0.5m/px) BEV space for an LR and HR BEV space, for X and Y axes, commonly. 1) A two-level FPN layer with 2-stride CNN and up-sampling layers, 2) a range crop operation, 3) and 4-scale SR are applied on the LR BEV. Thus, the scope manipulation procedure of ×$\times 4$ TNN is summarized as:
  • FPN layer: 2.0m/px $\mapsto$ 0.5m/px $\mapsto$ 2.0m/px
  • Range Crop: (--52.0m, 52.0m) $\mapsto$ (--50m, 50m)
  • ×$\times 4$ up-sampling: 2.0m/px $\mapsto$ 0.5m/px
  • Output: (-50m, 50m, 0.5m/px)
IV-D Training Strategy
Our optimization aims to implement a modular TNN mechanism. We first train a base model, which utilizes LR BEV. Afterward, we freeze the parameters before local restoration and fine-tune local restoration networks and decoder to use HR BEV space. This strategy enables attaching varying-scaled learnable up-sampling to the latter of the model. Furthermore, the designs address the concern that high-resolution BEV consumes large memory space generating partial backpropagation memory.
IV-E Implementation Details
Our implementations are built on top of mmdetections [5], [6]. We use the Swin-T (Tiny) [17] model and feature lifting module [21] to encode the camera BEV. We adopt sparse convolution [16] as a LiDAR backbone. To fuse the LiDAR and the camera BEV features, we use a fully convolutional layer with Feature Pyramid Networks (FPN) [14]. The data augmentations we used are the same methods as the baseline work [18]. Detailed model and training configurations are shown in our publicly opened codes.
V EXPERIMENTAL RESULTS
In our experiments, we concentrate on exploring the cost-efficiency of the TNN mechanism, its impact on local restoration for perception, and the compatibility of TNN with both LiDAR and camera in Tables I, II and III. Then we analyze model efficiency with Fig. 5(a) and Fig. 5(b).
Figure 7: HR perception using ×$\times 8$ TNN. Maps have 0.1m/px inner and 0.8m/px outer scopes.
V-A Quantitative Results
Setup. In Table I, we compare trainable up-samplings with the hand-crafted approaches: ‘Nearest’ neighbor, ‘Bilinear’, and ‘Bicubic’ interpolations, to validate the importance of the learnable local restoration. In the table, we examine the mean Intersection over Union (mIoU) alongside the inference time latency and the training time maximum memory of a GPU. The reported memory usage and latency are measured based on a single batch, and the memory excludes cache memory. In Table II, we validate the TNN compatibility to camera, LiDAR, and LiDAR-camera fusion modality cases. In both Table I and Table II, we use (-50m, 50m, 2.0m) BEV range and restore the extracted BEV features to (-50m, 50m, 0.5m) scope, applying a ×$\times 4$ TNN. In Table I, we compare scale factors of ×2,×4$\times$ 2$\times$ 4× 2 , × 4, and ×$\times 8$ to explore the representation power of LR BEV. In Table IV, we assess the descriptive capabilities of both LR and HR BEVs to analyze the difference of both spaces. As an evaluation metric, we use mIoU protocols of BEVFusion [18]. Gray shading in each table represents the specifications of our base model.
Local Restoration. The comparison in Table I reveals three insights. 1) Learnable up-samplings (Pixel shuffle, deconvolution) show superior BEV representation compared to handcrafted methods (Nearest neighbor, Bilinear, and Bicubic interpolations). This suggests that non-linear aliasing and blocky artifacts in BEV can be effectively restored using neural networks. 2) Deconvolution is an over-determined form of pixel shuffle as analyzed by Shi et al. [28]. Both methods yield similar performance although pixel shuffle is more cost-effective. 3) The TNN mechanism’s learnable up-sampling approach offers a reasonable trade-off between performance and costs. Compared to non-learnable methods, it achieves 14.2% mIoU gain with an additional 0.4GB GPU memory consumption and a 6ms latency increase in comparison to pixel shuffle over Bicubic interpolation. In Table IV, we analyze the restoration performance. For this purpose, we establish a baseline (or BEVFusion) trained under two different conditions: a 0.5m/px HR BEV space and a 2.0m/px LR BEV space. Additionally, we include a base model equipped with our proposed method for comparison. TNN uses 2.0m/px for LR BEV and up-samples it to 0.5m/px BEV resolution. As shown in the table, BEV map construction in LR space typically offers superior perception for understanding large urban scenes due to the standardized shapes. Our mechanism leverages this property by locally restoring the up-sampled BEV representation.
Modality Compatibility. In Table II, we evaluate the compatibility of the TNN mechanism across various modalities, including LiDAR, camera, and LiDAR-camera fusion. As indicated in the table, all modalities exhibit performance gains, which can be attributed to the superior representation power of LR BEV for semantic map construction, coupled with local restoration to evaluate HR BEV.
@@ V-A Quantitative Results
Representation Power. To delve into a trade-off between efficiency and representation power, we investigate the map construction performances of varying-scaled TNN in Table III with our LiDAR-camera fusion baseline, BEVFusion [18]. The exploration assesses mIoU, computation latency, and maximum GPU memory consumption. ×$\times 2$, ×$\times 4$, and ×$\times 8$ scale models are respectively trained under 0.5m/px, 1.0m/px, 2.0m/px, and 4.0m/px BEV resolutions for fair comparison in (50m, 50m, 0.5m/px) BEV space. The results imply there is the most reasonable scale factor to represent BEV providing powerful perception with proper memory consumption as the results of ×$\times 4$ scale. We see that the memory consumption of ×$\times 2$ requires the largest memories due to the additional memory of learnable up-sampling.
@@
Efficiency. In Fig. 5(b) and Fig. 5(a), we show the efficiency of our mechanism across increasingly heavier architectural configurations, illustrating how TNNs maintain efficiency even as the system architecture grows. Specifically, we add MSA layers, a well-known operator requiring large memory space, to the fusion neck of BEVFusion and TNN. In Fig. 5(a), BEVFusion exhibits increasing latency, and Fig. 5(b) shows diverging memory. In contrast, our model maintains stable cost consumption even when incorporating 32 MSA layers into its architecture. This property will contribute to applying heavy training tricks like vision foundation models.
@@ V-B Qualitative Results
Local Restoration. In Fig. 3, we compare up-sampling methods qualitatively to demonstrate the limited map construction of the hand-crafted approaches (nearest neighbor, bilinear, bicubic interpolations). The images of the first row are camera images, and the first column of the second row shows LiDAR points of BEV. The hand-crafted methods fail to reconstruct blocky LR BEV showing uncompleted construction of deteriorated class labels like walkways, and road lanes. In addition, they fail to restore aliasing as in the edges of “walkway” labels. However, our pixel shuffle-based learnable upsampler resolves the problems showing successful restorations of the aliased and blocky BEV features.
Up-scaling Factor. In Fig. 4, we explore various up-scale factors, qualitatively. TNNs with low up-scale factors (×1,×2$\times$ 1$\times$ 2× 1 , × 2) fail to capture high-level scene features but preserve local details as shown in the black and green ROIs, respectively. In contrast, TNNs with high up-scale factors (×4,×8$\times$ 4$\times$ 8× 4 , × 8) facilitate approximating the high-level description. However, excessive TNN with up-scale factors like ×$\times 8$ cause severe signal aliasing and blocky artifacts, thereby imposing a reasonable selection of a specific point that has compactness and tolerant signal quality.
@@
HR BEV Perception. In Fig. 7, we demonstrate our mechanism’s HR perception. The task severely limits BEVFusion to training its model due to the slow training speed and huge backpropagation memory. However, the model with our mechanism is free from the diverging training costs of learning HR BEV representation as our ×$\times 8$ TNN model, which respectively adopts 0.8m/px and 0.1m/px for LR and HR BEV resolution, has efficient training pipelines. Our proposed mechanism enables efficient training, facilitates HR perception, and enhances map construction performance.
@@ VI CONCLUSIONS
We suggested the trumpet neural network (TNN), a novel architecture tailored with a local restoration focus, aimed at the precise construction of bird’s eye view (BEV) maps. TNN resolved diverging training costs, thereby facilitating the training for the extraction of high-resolution BEV maps. Our experiments in semantic BEV map construction demonstrated that our proposed approach finely, and precisely restores BEV, achieving 8.2% mIoU gain over the existing baseline model, BEVFusion. Moreover, the adaptability and effectiveness of the TNN mechanism are validated across various sensory modalities, including camera, LiDAR, and LiDAR-camera fusion, respectively showing 4.3%, 10.1%, and 8.2% mIoU gains compared to the same models without the TNN mechanism. The experimental results show the solid compatibility and effectiveness of TNN, establishing a new baseline for precise BEV map construction which is essential for safe autonomous driving.

### v4

@@ V-A Quantitative Results
We incrementally add an MSA layer to the fusion neck and compare their costs. X and Y axes denote costs and the number of MSA layers, respectively.
Figure 4: Comparison of BEVRestore Restoration to conventional methods. We use a (-50m, 50m) BEV scope with 2.0m/px LR BEV and 0.5m/px HR BEV resolutions. The hand-crafted Nearest, Bilinear, and Bicubic methods suffer from the incomplete restoration of aliasing and blocky artifacts, compared to the BEVRestore, which utilizes learnable restoration.
@@
Representation Power. To delve into a trade-off between efficiency and representation power, we investigate the map construction performances of varying-scaled BEVRestore in Table V with our LiDAR-camera fusion baseline, BEVFusion [19]. The exploration assesses mIoU, computation latency, and maximum GPU memory consumption. We restore 0.5m/px, 1.0m/px, 2.0m/px, and 4.0m/px LR BEVs to 0.5m/px for fair comparison in (50m, 50m, 0.5m/px) BEV space. As in the table, the results imply there is a reasonable scale factor for a powerful representation of BEV with proper memory consumption as the results of the ×$\times 4$ scale. Although our mechanism requires additional memory to restore BEV, as shown in Fig. 3(b), it defenses any models across sensor modalities and perception tasks from diverging training costs caused by computationally intensive modules, such as deep nets with increased parameters and attention layers [30].
Table VI: Comparisons on cost footprints.
(a) (50.0m, 50.0m, 0.5m/px), BEVFusion [19].
Measure Coverage Mem. Latency
BEV Encoder Neck - Decoder Cache (GB) (ms)
2.3 158
2.4 167
2.6 176
3.6 176
(b) (50.0m, 50.0m, 0.5m/px), Enhancement.
Measure Coverage Mem. Latency
BEV Encoder Neck Restore Decoder Cache (GB) (ms)
2.3 150
2.4 161
4.0 207
4.1 218
5.1 218
(c) (50.0m, 50.0m, 1.0m/px $\rightarrow$ 0.5m/px).
Measure Coverage Mem. Latency
BEV Encoder Neck Restore Decoder Cache (GB) (ms)
2.2 146
2.3 155
2.9 176
3.1 187
4.2 187
(d) (50.0m, 50.0m, 2.0m/px $\rightarrow$ 0.5m/px).
Measure Coverage Mem. Latency
BEV Encoder Neck Restore Decoder Cache (GB) (ms)
2.1 134
2.1 142
2.5 155
2.7 164
3.8 164
(e) (50.0m, 50.0m, 4.0m/px $\rightarrow$ 0.5m/px).
Measure Coverage Mem. Latency
BEV Encoder Neck Restore Decoder Cache (GB) (ms)
2.1 129
2.1 133
2.4 144
2.6 153
3.7 153
@@
Efficiency. In Fig. 3(a) and Fig. 3(b), we show the efficiency of our mechanism across increasingly heavier architectural configurations, illustrating how BEVRestore maintains efficiency even as the system architecture grows. Specifically, we add multi-head self-attention (MSA) layers, a well-known operator requiring large memory space, to the fusion neck of BEVFusion and BEVRestore. In Fig. 3(a), BEVFusion exhibits increasing latency, and Fig. 3(b) shows diverging memory. In contrast, our model maintains stable cost consumption even when incorporating 32 MSA layers into its architecture. This property will contribute to applying heavy training tricks like vision foundation models. In Table VI, we investigate the cost footprints of our BEVRestore models and baseline. “Cache” denotes cache memory of tensors excluding models’ consumption. “BEV Encoder” consists of modality backbones, encoder FPN, and a view transformer, “Neck” includes fusion and the FPN of the fused BEV feature, “Restore” is our BEVRestore network, and “Decoder” means CNN networks for decoding into semantic maps. As shown in the table, although our proposed mechanism requires additional memory for BEVRestore, its utilization of LR BEV contributes to the low cost of the BEV encoder and the neck module’s BEV enhancement.
Figure 5: Restoration Comparisons on varying up-sampling factors. We use a (-50m, 50m) BEV scope with 0.5m/px, 1.0m/px, 2.0m/px, 4.0m/px LR BEV resolutions for ×1,×2,×4$\times$ 1$\times$ 2$\times$ 4× 1 , × 2 , × 4, and ×$\times 8$ models, respectively. Their HR BEV space utilizes 0.5m/px HR resolution. The ×$\times 1$ and ×$\times 2$ BEVRestores fail to learn high-level scene features (Black ROIs). In contrast, ×$\times 8$ BEVRestore suffers from inaccurate prediction (Green ROIs) caused by severe loss of prior, compared to the ×$\times 4$ BEVRestore.
Figure 6: HR perception using ×$\times 8$ BEVRestore. Maps have 0.1m/px inner and 0.8m/px outer scopes.
Figure 7: Overconfident Issues. Green ROIs: Far regions occluded by cars. Our map construction struggles with the overconfident prediction of uncertain regions.
@@ V-B Qualitative Results
BEVRestore. In Fig. 4, we compare up-sampling methods qualitatively to demonstrate the limited map construction of the hand-crafted approaches (nearest neighbor, bilinear, bicubic interpolations). The images of the first row are camera images, and the first column of the second row shows LiDAR points of BEV. The hand-crafted methods fail to reconstruct blocky LR BEV showing uncompleted construction of deteriorated class labels like walkways, and road lanes. In addition, they fail to restore aliasing as in the edges of “walkway” labels. However, our pixel shuffle-based learnable upsampler resolves the problems showing successful restorations of the aliased and blocky BEV features.
Up-scaling Factor. In Fig. 5, we explore various up-scale factors, qualitatively. BEVRestores with low up-scale factors (×1,×2$\times$ 1$\times$ 2× 1 , × 2) fail to capture high-level scene features but preserve local details as shown in the black and green ROIs, respectively. In contrast, BEVRestores with high up-scale factors (×4,×8$\times$ 4$\times$ 8× 4 , × 8) facilitate approximating the high-level description. However, excessive BEVRestore with up-scale factors like ×$\times 8$ cause severe signal aliasing and blocky artifacts, thereby imposing a reasonable selection of a specific point that has compactness and tolerant signal quality.
HR BEV Perception. In Fig. 6, we demonstrate our mechanism’s HR perception. The task severely limits BEVFusion to training its model due to the slow training speed and huge backpropagation memory. However, the model with our mechanism is free from the diverging training costs of learning HR BEV representation as our ×$\times 8$ BEVRestore model, which respectively adopts 0.8m/px and 0.1m/px for LR and HR BEV resolution, has efficient training pipelines. Our proposed mechanism enables efficient training, facilitates HR perception, and enhances map construction performance.
@@ VI CONCLUSIONS
We suggested the BEVRestore, a novel mechanism tailored with a BEVRestore, aimed at a memory-efficient precise construction of bird’s eye view (BEV) maps. BEVRestore resolved diverging training costs, thereby facilitating the training for the extraction of high-resolution BEV features. Our experiments in BEV map construction demonstrated that our proposed approach finely, and precisely restores BEV features, achieving 4.3%, 10.1%, and 8.2% mIoU gains on BEV segmentation with camera, LiDAR, and LiDAR-camera fusion, respectively, over the existing baseline model BEVFusion. Moreover, the adaptability and effectiveness of the BEVRestore mechanism for vectorized map construction are validated respectively showing the mIoU gains of 9.1% mAP and 2.2% mIoU for vectorized and rasterized map construction of HDMapNet and 1.2% mIoU gain for rasterized map construction of BEVFormer. The experimental results show the solid compatibility and memory efficiency of BEVRestore, establishing a new baseline for precise BEV map construction which is essential for safe autonomous driving.